OpenStudy (anonymous):

solve the system using the subtitution method 3x+6y=3 x-2y=5

hero (hero):

Solve for y: \[y = \frac{3-3x}{6}\] \[y=-\frac{5-x}{2}\] Substitute y = y: \[\frac{3-3x}{6} = -\frac{5-x}{2}\] Cross Multiply: \[6(5-2x) = -2(3-3x) \] Simplify: \[30 -12x = -6 +6x \]\[36 = 18x \]\[x = 2 \]\[y = 1/2\]

hero (hero):

Any questions?